OpenStudy (sleepyjess):

\(\bf sin15\) \(\dfrac{\sqrt{6}+\sqrt{2}}{4}\) \(\dfrac{\sqrt{6}-\sqrt{2}}{4}\) \(-\dfrac{\sqrt{6}-\sqrt{2}}{4}\) \(-\dfrac{\sqrt{6}+\sqrt{2}}{4}\)

OpenStudy (sleepyjess):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

15 = 45 - 30 use sin(x-y) = sin(x)cos(y) - cos(x)sin(y)

jimthompson5910 (jim_thompson5910):

x = 45 y = 30

OpenStudy (sleepyjess):

I am thinking the second one

jimthompson5910 (jim_thompson5910):

yeah I'm getting \[\Large \frac{\sqrt{6}-\sqrt{2}}{4}\] as well